OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

If it is parallel then the equation must be y + t where t is a real number and y represents the equation given, because that will just go up with the line or down, making it parallel. Then you substitute the point you want: x=-2, wich gives y=7, but you want it to be -3, so t must be -10. So the equation is y'=x + 9 - 10 = x-1

OpenStudy (anonymous):

the answers to choose from are x – y = 1 x + y = 1 x – y = –5 x + y = –5

OpenStudy (anonymous):

You just need to change the equations, my answer was y = x - 1, then it would be the first option.

OpenStudy (anonymous):

okay, thank you!